# Diffusion in Zero-Shot Learning for Environmental Audio

### v1

@@ VI Discussion
In a true zero-shot learning real-world application, it is hard to know how well a model will perform on the target class due to a lack of available samples. Therefore, consistency across training runs is important. While ZeroDiffusion has a superior accuracy-based performance, the bilinear method has superior consistency. This is an important consideration when choosing an appropriate method for zero-shot learning in practical applications.
The hubness problem is an issue when considering practical usage because of the poor accuracy on specific individual classes that are overshadowed by the hubs. Many of these specific classes are predicted at a rate worse than a random model. Hubness appears to correlate to overall model performance, as seen when comparing the confusion matrices and overall accuracy of two partitions of the ESC-50 dataset.
One key avenue in reducing the hubness problem is first understanding why the different partitions used in this work had varying levels of hubness. Partitions can be crafted for this purpose to identify the effect of different criteria on hubness. One question is whether the unseen classes, seen classes, or both, influence the performance. If particular characteristics in the seen class set influence the hubness problem, then the seen class set can be supplemented with classes exhibiting these characteristics. A more likely scenario is that both influence the performance, and the seen classes used in training may need to be curated for the characteristics of the specific unseen set in the given problem.
Zero-shot learning lacks theoretical foundations, with methods developing through empirical results. There is little understanding of upper bounds to accuracy for a given unseen and seen class sets. A given seen class set will not necessarily have enough information to extrapolate to particular unseen classes. A trivial example is if the seen class set is the empty set. A deeper understanding of these requirements may be an avenue to reducing hubness, and therefore improving overall classification performance.
@@ VII Conclusion
A comparison of three generative methods and the baseline method in environmental audio zero-shot learning has been presented. Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly or worse than the baseline, and present higher variance.
The third generative method, ZeroDiffusion, is a novel diffusion-based method for zero-shot learning. It shows a higher accuracy performance than the other three methods, thus producing a new state-of-the-art result in the domain of environmental audio on ESC-50 and FSC22. Diffusion shows potential in zero-shot learning to improve the performance across domains. Future work could include benchmarks for diffusion-based zero-shot learning models on datasets in areas other than environmental audio.
An analysis of confusion matrices and t-SNE graphs of the audio embeddings reveals the impact of audio embedding similarities on the hubness problem and model performance. This analysis is valuable since these similarities may not be intuitive from the raw audio data.

### v2

@@ VI Discussion
Given the results of this four-model and six-dataset study, three areas are considered: model consistency and the importance of consistency in zero-shot learning; the impact of dataset size; and the issue of theoretical guarantees in zero-shot learning.
VI-A Consistency of Methods
In a true zero-shot learning real-world application, it is hard to know how well a model will perform on the target class due to a lack of available samples. Therefore, consistency across training runs and datasets is important for building trust in the model. The issue of the black-box nature of machine learning models hinders application to key global problems, and the added uncertainty of model performance further exasperates these problems.
When considering the standard deviation of each result in Table II, there is a high variance trend amongst the generative methods (LisGAN, CADA-VAE, ZeroDiffusion), with CADA-VAE reaching a high 8.08% standard deviation on the UrbanSound8k dataset and ZeroDiffusion not far behind with a 7.77% deviation. Meanwhile, the highest standard deviation for ALE is around 1% on ESC-50 partitions and GTZAN. This is a serious problem for some domains and indicates a need to develop more stable generative methods to make use of the improved accuracy of these methods.
Between datasets, there is also an inconsistency, as noted with ESC-50 having top results across partitions from every method used in the presented experiments. ZeroDiffusion has high accuracy on ARCA23K-FSD; however, it suffers issues with fold 3 and the test fold despite large improvements on the other folds. Prediction of model performance across partitions and datasets can better inform the application and improvement of these methods.
VI-A1 Data Diversity
From the results presented, larger datasets typically faired better with ZeroDiffusion. TAU 2019 may have only ten classes, but with 1,440 samples per class it represents a large audio dataset and this appears to result in an accuracy improvement over other methods. The smaller two datasets, ESC-50 (40 samples per class) and FSC22 (75 samples per class) experienced instability in performance across methods. ARCA23K-FSD provided a balance between the number of classes (70) and the number of samples per class (on average 339 samples), and the methods were more consistent in their performance, with ZeroDiffusion dominating. The dataset size, both in the number of classes and samples per class, may influence which zero-shot classification method is more suitable for individual scenarios and tasks. This is an important consideration when applying zero-shot learning in practice.
VI-A2 Theoretical Guarantees
Zero-shot learning lacks theoretical foundations, with methods developing through empirical results, as done in the presented work. There is little understanding of upper bounds on accuracy for a given set of unseen and seen classes. A given seen class set will not necessarily have enough information to extrapolate to particular unseen classes. A trivial example is if the seen class set is the empty set, then there is no information to learn. One indication of these theoretical guarantees is the differences between dataset sizes in the presented experiments and the effect on model stability. A deeper understanding of these requirements may improve overall classification performance.
@@ VII Conclusion
This study presents a comparison of three generative methods and a baseline method in zero-shot environmental sound classification. Two established generative methods in the domain of computer vision, CADA-VAE and LisGAN, are shown to perform similarly to the baseline and present higher variance.
The third generative method, ZeroDiffusion, is our novel diffusion-based method for zero-shot learning. It has a higher accuracy performance than the other three methods in most cases on the six chosen datasets, with a new state-of-the-art result on ARCA23K-FSD and TAU Urban Acoustics 2019, and reasonable results on ESC-50, FSC22, UrbanSound8k and GTZAN. The inclusion of TAU 2019 and GTZAN shows that the method can perform in applications other than environmental sound classification, with acoustic scene classification and music genre classification represented by these datasets. Future work on diffusion-based zero-shot models would benefit from considering other domains, such as image classification.